After 8 years jone will be 5 times as old as he was 8 years ago. Find his present age.
Answer by stanbon(75887)

After 8 years jone will be 5 times as old as he was 8 years ago.
Find his present age.
-----
Equation:
j+8 = 5(j-8)
j+8 = 5j-40
4j = 48
j = 12 yrs (Jone's present age)
-------------
Cheers,
Stan H.
----------